Talamoto Co. manufactures a single product that goes through two processes – mixing and cooking. The following data pertainsMultiple Choice $2.88. $2.40. $4.80. $3.14. $5.42.
Answer #1

Unit transferred out = 33000+100000-22000 = 111000

Equivalent unit of Conversion = 111000+(22000*50%) = 122000

Total Cost of conversion = 167000+494200 = 661200

Cost per equivalent unit of conversion = 661200/122000 = 5.42

So answer is e) $5.42